OpenStudy (anonymous):

The graph of a square root function f is shown below. Find each of the following, if possible. a. f(11) b. f(-1)

OpenStudy (e.mccormick):

Know what it means when you put a number into a function? \(f(x)=\sqrt{x}\) \(f(11)=\sqrt{11}\) \(f(-1)=\sqrt{-1}\)

OpenStudy (anonymous):

Thats what I thought but then I looked at the back of the book that gives the answers and it was a. 3 b. Undefined

OpenStudy (e.mccormick):

Yes. You need to try and get the roots. They seem to be doing it from a graph, so it is an estimate.

OpenStudy (anonymous):

Im so sorry but what do you mean they are doing it from a graph there is no graph given to me. Am I suppose to create it

OpenStudy (e.mccormick):

It says "The graph of a square root function f is shown below."

OpenStudy (anonymous):

Also how do you get the roots for f(11) to 3

OpenStudy (e.mccormick):

Well, here is the graph. When x is 11, what is y closest to? https://www.desmos.com/calculator/rbjpplht6a

OpenStudy (anonymous):

OHHH OK THANK YOU SO MUCH!

OpenStudy (anonymous):

That helps way more!

OpenStudy (e.mccormick):

Yah, they are taling about this: https://www.desmos.com/calculator/14sws5aoxx I added the x lines for 11 and -1.

OpenStudy (anonymous):

Thank you for clearing that up :)

OpenStudy (e.mccormick):

As you can see, -1 never touches the square root graph. That is the undefined. The other one they just want to know where it hits near. The real root of 11 is closer to 3.32, but they are rounding to a whole number.

OpenStudy (anonymous):

yea which is why its undefined awesome!